DIRC Operations Weekly Report

Period: Saturday 2006/08/05 - Friday 2006/08/11

Oncall: Rolf Andreassen, Bradley Wogsland
Fast Monitoring: Bradley Wogsland, Rolf Andreassen

File location: $BFROOT/www/Detector/DIRC/WeeklyReports/report_20060813.html
Plot directory: $BFROOT/www/Detector/DIRC/WeeklyReports/Plots/report_20060813 (empty)

Contents


BaBar-PEP news

PEP peak lumi: 11.87*1033 (previous week: 10.8*1033)

BaBar integrated lumi: 4.56 fb-1 (previous week: 5.24 fb-1)

Run 5b PEP delivery: 4.73 fb-1

Run 5 total delivery: 148.86 fb-1 Run 5 total BaBar recorded: 143.70 fb-1 including ~5 fb-1 off-peak data

PEP activities this week:

PEP activities planned next week :

see run 5b plan

Foreseen work opportunities (PEP down periods):

see run 5b plan

BaBar activities/issues this week:

DIRC activities

Paged around suppertime August 8th about the water in the tank on the DQM checklist being around 1 cm too high. Assured the pilot that we were aware of this and altered the short term instructions to reflect this knowledge.

I would have been paged August 11th about high bin 11 (data truncation) by the swing DQM shifter, but it was me. Basically this was due to the high luminosity (11+). MCC was able to tune it down somewhat, getting it just under .1%, and since the data quality was unaffected (looking at the fast monitoring) Emmanuele said we could live with it. Watching the DIRC background over the next few days however, MCC managed to keep it way down while running at >11*1033.

See DIRC logbook for details

Backgrounds

Background summary plot

Scheduled activities

None

Pending Work

None

Manpower situation

Bradley took over on-call commissioning duties from Rolf August 8th.

Links